OpenStudy (anonymous):

It takes a smaller hose 3 times as long to fill a small swimming pool as it does a larger hose. It takes both hoses working together 30 minutes to fill the swimming pool. How long will it take the smaller hose to fill the pool by itself?

OpenStudy (baru):

let the larger hose take x min to fill the tank then the smaller hose takes 3x min let volume of the tank be V so rate of small pipe =V/3x galons/minute rate of large pipe=V/x gallons/minute rate together=v/3x + v/x volume filled by the pipes in 30 min is V (whole tank) thus 30(v/3x+v/x)=v

OpenStudy (baru):

\(30 \times( \frac{v}{3x} + \frac{v}{x})=v\)

OpenStudy (baru):

solve that, 'v' gets cancelled, it becomes an equation in one variable 'x'

OpenStudy (anonymous):

Let a and b be the smaller and larger fill times respectively. Solve the following for a and b:\[\left\{\frac{1}{a}+\frac{1}{b}=\frac{1}{30},\frac{b}{a}=3\right\} \]where 1/a means 1 filling per "a" minutes and so forth. a = 40 minutes and b = 120 minutes.

OpenStudy (lochana):

@niz1995 did you understand it?

OpenStudy (lochana):

this is what I think about your question, when both hoses fill the tank together, after 30 minutes, "1/4 of tank is filled by small tank" 

OpenStudy (lochana):

so if small hose takes 30 minutes to fill 1/4th of the tank, It will take 4*30 minutes for the small hose to fill whole tank answer is 120min